# Fairshare Data Pricing via Data Valuation for Large Language Models

### v2

@@ E.1 Data Valuation Experiments
Model Training: After obtaining purchasing decisions for all data samples, the buyers train their models using the purchased data. In order to conduct a fair comparison across buyers, we sample a set number of data from the buyers’ purchases (shown in Figure 6). We train each model (i.e., buyer) on these samples separately using LoRA [82] for 3 epochs, with a learning rate of 2e-5 and batch size 32. All models are trained on A6000 GPUs on single GPU settings and take less than 1 hour.
Model Evaluation: For evaluation, we use the test splits of the previously mentioned datasets. In particular, we use 5-shot evaluation on the MathQA test set, and 4-shot evaluation in on the MedQA test. Table 2 in Appendix F shows the demonstrations used for 5-shot and 4-shot evaluation.
Market/Pricing Setup: We reserve $1\%$ of the samples from each dataset’s training split to represent the existing data in their respective markets. Each data sample was randomly priced between $(0,1]$. Next, for each remaining data sample in the training set, we determine whether each buyer will purchase the data sample at potential price points $[0.5,0.625,0.75,0.875,1.0]$ by solving Equation 2. The seller then sets their prices according to Equation 4. We price data separately for each data valuation method. This assesses the method’s ability to discern whether a new data sample is worth purchasing for each buyer given the existing market data, as noted in our analysis in Section 3.
@@ E.2 Data Pricing Experiments
Robustness Check. As discussed in Lemma 7, the threshold $t^{*}$ when fairshare becomes optimal for the buyer) increases as the discount factor $\delta$ decreases. To run a robustness check, for the high-budget buyer in Figure 4(a), setting $\delta=0.999,0.99,0.98$ yields $t^{*}=32,38,44$ respectively, which is consistent with our theoretical analysis. In below, we compare the change of buyer’s accumulative utilities over different values of $\delta$.
@@ F.1 Algorithms
8:              to decide whether to purchase $D_{j}$ at price $p_{j}$triangleright See Eqn. 2
@@
10:         Seller computes net profit $r(p_{j})$ assuming price $p_{j}$triangleright See Eqn. 3
@@
13:     and sets $p_{j}^{*}$ as the fixed price for $D_{j}$triangleright See Eqn. 4
@@ F.2 Data Valuation Oracle Experiments
Influence-based methods, such as $\text{Infl}_{\text{IP}}$ approximate the influence of a sample $d$ on a $d^{\prime}$ for a model parameterized by $\theta$ by estimating the effects of training or “upweighting” (alternatively, removing/“downweighting”) on $d$ (see Appendix A). In past literature, $\text{Infl}_{\text{IP}}$ is validation through an “Oracle” (One-Step Training) score, which we denote as $\text{Oracle}(d,d^{\prime})=\mathcal{L}(d^{\prime};\theta)-\mathcal{L}(d^{\prime};\hat{\theta})$, where $\hat{\theta}=\theta-\eta\nabla\mathcal{L}(d;\theta)$ and $\eta$ is the learning rate [56, 62, 23].
To compare the difference between $\text{Infl}_{\text{IP}}$ valuation versus its oracle valuation, we conduct the same experiments described in Section 4.3. Figure 8 shows that $\text{Infl}_{\text{IP}}$ and Oracle have decent correlation in their agreement in their valuation of the sellers’ data, which supports findings in previous works [62]. We note that in the case where correlation is decent, such as in Figure 8(b), the final model performance between these methods is close, as seen in Figure 8(d). In the case where correlation is lower, such as in Figure 8(a), the final model performance between these methods initially have a gap, but become more similar as the amount of data purchased increases, as seen in Figure 8(c). This suggests that in practice, even in cases when the agreement between $\text{Infl}_{\text{IP}}$ and Oracle may not be very high, final model performance resulting from these two methods can still be similar.
@@ F.3 Additional Experimental Results
Figure 9: Buyers’ model (top row: Pythia-1b, bottom row: Pythia-410m) performance and costs from their purchased data from math, medical, and physical reasoning data markets. Purchasing decisions were using the constant, random, BM25, $\text{Infl}_{\text{IP}}$ data valuation methods (see Section 4.3 for details).
Figure 10: Analysis of (1) buyer’s cumulative utilities with high-budget buyer (Figures 10(a), 10(e) and 10(i)) and low-budget buyer (Figures 10(b), 10(f) and 10(j)), and (2) sellers’ average cumulative profits (Figures 10(c), 10(g) and 10(k)) and number of sellers in the market (Figures 10(d), 10(h) and 10(l)) over time ($T=100$). Model: Pythia-1b, Pythia-410m, and Llama-3.2-Inst.-1b; Task: medqaQA. Experimental groups: (1) fairshare, (2) reduced, (3) random, and (4) current pricing.
Figure 11: Analysis of (1) buyer’s cumulative utilities with high-budget buyer (Figures 11(a), 11(e) and 11(i)) and low-budget buyer (Figures 11(b), 11(f) and 11(j)), and (2) sellers’ average cumulative profits (Figures 11(c), 11(g) and 11(k)) and number of sellers in the market (Figures 11(d), 11(h) and 11(l)) over time ($T=100$). Model: Pythia-1b, Pythia-410m, and Llama-3.2-Inst.-1b; Task: MathQA. Experimental groups: (1) fairshare, (2) reduced, (3) random, and (4) current pricing.
Figure 12: Analysis of (1) buyer’s cumulative utilities with high-budget buyer (Figures 12(a), 12(e) and 12(i)) and low-budget buyer (Figures 12(b), 12(f) and 12(j)), and (2) sellers’ average cumulative profits (Figures 12(c), 12(g) and 12(k)) and number of sellers in the market (Figures 12(d), 12(h) and 12(l)) over time ($T=100$). Model: Pythia-1b, Pythia-410m, and Llama-3.2-Inst.-1b; Task: PIQA. Experimental groups: (1) fairshare, (2) reduced, (3) random, and (4) current pricing.
@@ 8. Answer: [Yes]
  9. Justification: We discuss our limitations in Appendix G
@@ 38. Answer: [Yes]
  39. Justification: We describe this in Section E.1
@@ 48. Answer: [Yes]
  49. Justification: We discuss this in Appendix G.

### v3

@@ E.1 Data Valuation Experiments
Model Training: After obtaining purchasing decisions for all data samples, the buyers train their models using the purchased data. In order to conduct a fair comparison across buyers, we sample a set number of data from the buyers’ purchases (shown in Figure˜6). We train each model (i.e., buyer) on these samples separately using LoRA [91] for 3 epochs, with a learning rate of 2e-5 and batch size 32. All models are trained on A6000 GPUs on single GPU settings and take less than 1 hour.
Model Evaluation: For evaluation, we use the test splits of the previously mentioned datasets. In particular, we use 5-shot evaluation on the MathQA test set, and 4-shot evaluation in on the MedQA test. Table˜2 in Appendix˜F shows the demonstrations used for 5-shot and 4-shot evaluation.
@@ E.2 Data Pricing Experiments
Robustness Check. As discussed in Lemma˜4, the threshold $t^{*}$ when fairshare becomes optimal for the buyer) increases as the discount factor $\delta$ decreases. To run a robustness check, for the high-budget buyer in Figure˜4(a), setting $\delta=0.999,0.99,0.98$ yields $t^{*}=32,38,44$ respectively, which is consistent with our theoretical analysis. In below, we compare the change of buyer’s accumulative utilities over different values of $\delta$.
@@ F.1 Algorithms
8:     to decide whether to purchase $D_{j}$ at price $p_{j}$ $\triangleright$ See Eqn. 2
@@
10:   Seller computes net profit $r(p_{j})$ assuming price $p_{j}$ $\triangleright$ See Eqn. 3
@@
13:  and sets $p_{j}^{*}$ as the fixed price for $D_{j}$ $\triangleright$ See Eqn. 4
@@ F.2 Data Valuation Oracle Experiments
[…]
To compare the difference between $\text{Infl}_{\text{IP}}$ valuation versus its oracle valuation, we conduct the same experiments described in Section˜4.3. Figure˜8 shows that $\text{Infl}_{\text{IP}}$ and Oracle have decent correlation in their agreement in their valuation of the sellers’ data, which supports findings in previous works [71]. We note that in the case where correlation is decent, such as in Figure˜8(b), the final model performance between these methods is close, as seen in Figure˜8(d). In the case where correlation is lower, such as in Figure˜8(a), the final model performance between these methods initially have a gap, but become more similar as the amount of data purchased increases, as seen in Figure˜8(c). This suggests that in practice, even in cases when the agreement between $\text{Infl}_{\text{IP}}$ and Oracle may not be very high, final model performance resulting from these two methods can still be similar.
@@ F.3 Additional Experimental Results
Figure 9: Buyers’ model (top row: Pythia-1b, bottom row: Pythia-410m) performance and costs from their purchased data from math, medical, and physical reasoning data markets. Purchasing decisions were using the constant, random, BM25, $\text{Infl}_{\text{IP}}$ data valuation methods (see Section˜4.3 for details).
Figure 10: Analysis of (1) buyer’s cumulative utilities with high-budget buyer (Figures˜10(a), 10(e) and 10(i)) and low-budget buyer (Figures˜10(b), 10(f) and 10(j)), and (2) sellers’ average cumulative profits (Figures˜10(c), 10(g) and 10(k)) and number of sellers in the market (Figures˜10(d), 10(h) and 10(l)) over time ($T=100$). Model: Pythia-1b, Pythia-410m, and Llama-3.2-Inst.-1b; Task: medqaQA. Experimental groups: (1) fairshare, (2) reduced, (3) random, and (4) current pricing.
Figure 11: Analysis of (1) buyer’s cumulative utilities with high-budget buyer (Figures˜11(a), 11(e) and 11(i)) and low-budget buyer (Figures˜11(b), 11(f) and 11(j)), and (2) sellers’ average cumulative profits (Figures˜11(c), 11(g) and 11(k)) and number of sellers in the market (Figures˜11(d), 11(h) and 11(l)) over time ($T=100$). Model: Pythia-1b, Pythia-410m, and Llama-3.2-Inst.-1b; Task: MathQA. Experimental groups: (1) fairshare, (2) reduced, (3) random, and (4) current pricing.
Figure 12: Analysis of (1) buyer’s cumulative utilities with high-budget buyer (Figures˜12(a), 12(e) and 12(i)) and low-budget buyer (Figures˜12(b), 12(f) and 12(j)), and (2) sellers’ average cumulative profits (Figures˜12(c), 12(g) and 12(k)) and number of sellers in the market (Figures˜12(d), 12(h) and 12(l)) over time ($T=100$). Model: Pythia-1b, Pythia-410m, and Llama-3.2-Inst.-1b; Task: PIQA. Experimental groups: (1) fairshare, (2) reduced, (3) random, and (4) current pricing.
@@ 8. Answer: [Yes]
  9. Justification: We discuss our limitations in Appendix˜G
@@ 38. Answer: [Yes]
  39. Justification: We describe this in Section˜E.1
@@ 48. Answer: [Yes]
  49. Justification: We discuss this in Appendix˜G.